the sum of two consecutive even integers is 118.
what would this equation look like?

Answer by wilft1(217)

the sum of 2 consecutive integers is 118
x + x + 2 = 118
subtract 2 from both sides
x + x = 116
combine like terms
2x = 116
divide both sides by 2
x = 58
first integer is 58
x + 2 = 60
second integer is 60